OpenStudy (anonymous):

what is the slope of the line that passes through the points (-2 5) and (1 4) 1.)-3 2.)-2 3.)-1/3 4.)1/3

OpenStudy (anonymous):

m = slope = (y2-y1) / (x2-x1)

OpenStudy (anonymous):

In other words... (5-4) / (-2-1) = slope = m

OpenStudy (anonymous):

Nvm listen to RyleeC I derped with the numbers a bit

OpenStudy (anonymous):

@ohlookidontunderstandanotherthing

OpenStudy (anonymous):

You had it right doesn't matter which order, as long as your order pairs are consistent.

OpenStudy (anonymous):

Okay! Sorry! I'll let you handle this question!

OpenStudy (anonymous):

but (5-4) / (-2-1) isn't the numbers

OpenStudy (anonymous):

You have to simplify it...

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

1/-3 = -1/3

OpenStudy (anonymous):

5-4 = 1 -2-1 = -3 so 1/-3

OpenStudy (mayankdevnani):

@cody7744 http://openstudy.com/users/mayankdevnani#/updates/544bcbffe4b0e4617ebfc972 please go through this link

OpenStudy (anonymous):

so its 3

OpenStudy (mayankdevnani):

hope it helps !!!

OpenStudy (anonymous):

No, its -.333

OpenStudy (anonymous):

A fraction it is.

OpenStudy (anonymous):

so 4

OpenStudy (anonymous):

No. Read above and look at my formula.

OpenStudy (anonymous):

ok thxs

OpenStudy (mayankdevnani):

any problem while reading or observing the answer which is given by me :- http://openstudy.com/users/mayankdevnani#/updates/544bcbffe4b0e4617ebfc972

OpenStudy (mayankdevnani):

@cody7744

OpenStudy (mayankdevnani):

REMEMBER :- Use STRAIGHT FORWARD FORMULA

OpenStudy (mayankdevnani):

just plug all the values and get your answer :)

OpenStudy (mayankdevnani):

so can you show your work ???